Galileo Galilei conducted experiments involving Italy’s famous Leaning Tower of Pisa to investigate the relationship between an object’s speed of fall and its weight. The Leaning Tower is 804 ft. shorter than the Eiffel Tower in Paris, France. The two towers have a total height of 1,164 ft. How tall is each tower?
Found 2 solutions by Cromlix, stanbon:
Answer by Cromlix(4381)

Hi there,
Assume that the Eiffel Tower = 'x'
Leaning Tower = x - 804.
Therefore:
x + (x - 804) = 1,164 ft
x + x - 804 = 1,164 ft
2x - 804 = 1,164
2x = 1,164 + 804
2x = 1,968
x = 984 ft
Eiffel Tower = 984 ft
The Leaning Tower = 180 ft
Hope this helps :-)

Answer by stanbon(75887)

Galileo Galilei conducted experiments involving Italy’s famous Leaning Tower of Pisa to investigate the relationship between an object’s speed of fall and its weight. The Leaning Tower is 804 ft. shorter than the Eiffel Tower in Paris, France. The two towers have a total height of 1,164 ft. How tall is each tower?
----
LT = ET-804
LT + ET = 1164
------
ET-804 + ET = 1164
2*ET = 1908
ET = 954 ft (height of Eiffel Tower)
LT = 954-804 = 150 (height of Leaning Tower)
============
Cheers,
Stan H.
-------------